Assume the MPC = 0.80. And, assume that the economy suddenly moves away from the stability condition and a recessionary gap emerges. If this initial movement away from full-employment GDP originated via a $200 billion decrease in private sector business investment spending (I), AD will eventually _______ by a cumulative total of _______.
Answer #1

Answer
AD will eventually decrease by a cumulative total of $1000 billion.
---------

The investment spending decreases mean AD will decrease. the total effect of it is a
total effect on AD=change in investment *multiplier
Multiplier =1/(1-MPC)=1/(1-0.8)=5
change in AD=5*200=$1000 billion